OpenStudy (yanasidlinskiy):

The area of a rectangular garden is 54 square meters. The width is 7 meters longer than one-third of the length. Find the length and the width of the garden. Use the formula, are = length * width.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

The width is 7 meters longer than one-third the length, so we can say \[\Large W = \frac{1}{3}L + 7\]

jimthompson5910 (jim_thompson5910):

The area of the rectangle is length * width, so \[\Large A = L*W\] \[\Large 54 = L*\left(\frac{1}{3}L + 7\right)\] Do you see how to solve for L from here?

OpenStudy (yanasidlinskiy):

Yeah. Would it be \[54=\frac{ 1 }{ 3 }x^2+7x\] From there I need help...

jimthompson5910 (jim_thompson5910):

what I would do from there is multiply everything by 3 to get rid of the fraction

jimthompson5910 (jim_thompson5910):

then get everything to one side

OpenStudy (yanasidlinskiy):

So like multiply 1/3x^2(3) and multiply 7x(3)?

jimthompson5910 (jim_thompson5910):

yeah and 54 as well

OpenStudy (yanasidlinskiy):

Okie, one sec.

OpenStudy (yanasidlinskiy):

Got it. Would it be: \[162=1x^2+21x\]

OpenStudy (yanasidlinskiy):

Or instead I should just put \[x^2\]

jimthompson5910 (jim_thompson5910):

either works

jimthompson5910 (jim_thompson5910):

now move that 162 over

OpenStudy (yanasidlinskiy):

Ok. I have a question... Does it matter with what number I multiply the numbers with?

jimthompson5910 (jim_thompson5910):

you mean like how you multiplied everything by 3 just now?

OpenStudy (yanasidlinskiy):

Yeah.

jimthompson5910 (jim_thompson5910):

The denominator of the fraction was 3, so multiplying everything by that clears out the denominator (ie eliminates the fraction)

jimthompson5910 (jim_thompson5910):

if you picked another number that wasn't a multiple of 3, then the denominator or fraction wouldn't go away

OpenStudy (yanasidlinskiy):

Oh ok! Gotcha! Ok. So I got: 1x^2+21x-162=0 Give me a sec. I wanna do it myself and you tell me what I got wrong.

OpenStudy (yanasidlinskiy):

I got: (x+54)(x-3) PLEASE tell me it's right haha! I've had so much trouble with this today.

jimthompson5910 (jim_thompson5910):

54 times -3 = -162 .... true 54 plus -3 = 21 ... false so that's not the correct factorization

OpenStudy (yanasidlinskiy):

I'm confused and I think this is the stop. So how would I know which number goes where? Or more like why does it matter where I put down the sign with which number?

OpenStudy (yanasidlinskiy):

*spot

jimthompson5910 (jim_thompson5910):

The quadratic formula may be easier to work with. Did you want to use that?

OpenStudy (yanasidlinskiy):

Yeah, I wanna use that instead... This is really hard for me and I have no idea why...

OpenStudy (yanasidlinskiy):

Ok, let me figure it out using the quadratic formula.

jimthompson5910 (jim_thompson5910):

1x^2+21x-162 means a = 1 b = 21 c = -162 plug those values into \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (yanasidlinskiy):

I'm not exactly sure if this is right, but so far I have: \[x=\frac{ -21\pm \sqrt{207} }{ 2 }\] Is that right?

OpenStudy (yanasidlinskiy):

*-201

jimthompson5910 (jim_thompson5910):

The 207 portion is incorrect b^2 - 4ac = 21^2 - 4*1*(-162) = 1089

OpenStudy (yanasidlinskiy):

So I multiplied it incorrectly? How?

jimthompson5910 (jim_thompson5910):

21^2 = 441 4*1*(-162) = -648

jimthompson5910 (jim_thompson5910):

21^2 - 4*1*(-162) = 441 - (-648) = 1089

OpenStudy (yanasidlinskiy):

Ahh! That's what I was missing! Gotcha! Ok. Hang on:)

OpenStudy (yanasidlinskiy):

So it'd be like \[x=\frac{ -21\pm \sqrt{1089} }{ 2 }\] ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (yanasidlinskiy):

Sorry for making you wait.... I'm kinda slow on thinking right now...

jimthompson5910 (jim_thompson5910):

You're doing great. Don't worry.

OpenStudy (yanasidlinskiy):

I hope so... So my answer would be x=534?

OpenStudy (yanasidlinskiy):

The answer can never be negative.. I know that...

jimthompson5910 (jim_thompson5910):

no

OpenStudy (yanasidlinskiy):

Ughgh. Ok, I need some real help. What is it?

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{ -21\pm \sqrt{1089} }{ 2 }\] \[\Large x=\frac{ -21+ \sqrt{1089} }{ 2 } \ \text{ or } \ x=\frac{ -21- \sqrt{1089} }{ 2 }\] \[\Large x=\frac{ -21+ 33 }{ 2 } \ \text{ or } \ x=\frac{ -21- 33 }{ 2 }\] \[\Large x=?? \ \text{ or } \ x=??\]

OpenStudy (yanasidlinskiy):

Wow, ok. I divided 1089 by 2 so that's what threw me off... Hang on let me get the final answer..

OpenStudy (yanasidlinskiy):

I got 6, -27.

jimthompson5910 (jim_thompson5910):

x took the place of the length L so L = 6 or L = -27 but a negative length makes no sense which is why we ignore -27

OpenStudy (yanasidlinskiy):

So, what would be my width? I don't seem to be understanding that part..

jimthompson5910 (jim_thompson5910):

go back to \[\Large W = \frac{1}{3}L + 7\]

jimthompson5910 (jim_thompson5910):

plug in L = 6

jimthompson5910 (jim_thompson5910):

or you can ask yourself "something times 6 gives the area of 54, what is that something?"

OpenStudy (yanasidlinskiy):

Oh yeah. Sorry. I'm like dead tired and didn't get enough sleep lol.. 9

jimthompson5910 (jim_thompson5910):

yep L = 6 and W = 9

OpenStudy (yanasidlinskiy):

Anyways, THANK YOU SO MUCH!!!!!! I can't appreciate you enough. I'd be sitting behind the computer sending you billions of messages with different thank you's. Thanks again!

jimthompson5910 (jim_thompson5910):

I'm glad it's making sense now

OpenStudy (yanasidlinskiy):

Yeah, me too. I have another one similar to this, but I'll figure it out myself and let you check it since it's coming to my brain lol. Anyways, just expect a "check message" from me or something...

jimthompson5910 (jim_thompson5910):

ok sounds like a good plan to me